# Automated Static Vulnerability Detection via a Holistic Neuro-symbolic Approach

### v1

@@ 6.4. Ablation Study
We conduct a comprehensive ablation study to measure the individual contributions of the techniques on the testing dataset. We use the F1 score ($2\cdot\frac{Prec.\cdot Rec.}{Prec.+Rec.}$) as our evaluation metric, as it captures both precision and recall, and is well-suited for imbalanced data where non-vulnerable code is much more than vulnerable code. Due to the space limit, we put the efficiency analysis of query merging in Appendix D.
@@ 6.4.1. Vulnerability Examples
MoCQ takes as input a few vulnerability examples to generate queries. The queries might perform better when generated with more examples. Here, we gradually added more projects and applied the queries to the testing dataset. This incremental setup helps us understand the impact of additional projects on performance. We conduct our experiments at the project granularity rather than the vulnerability granularity due to similar data leakage concerns (discussed in § 6.1).
Figure 5 illustrates the F1 scores on the same testing dataset under different numbers of input projects. The F1 score increased as more projects were used for query generation. However, the point at which performance stabilized varies by vulnerability type. For example, JavaScript XSS achieved a stably high F1 score with four projects, whereas JavaScript prototype pollution generally required ten. Overall, we find that ten projects are typically sufficient to generate high-quality queries. This corresponds to the 5-18 vulnerability examples across different vulnerability types (details in Appendix B).
Figure 5. F1 scores with queries generated with different numbers of input projects.
@@ 6.4.2. DSL Subset
To understand the effect of the subsetting technique, we developed a variant containing the full set of DSL, denoted as DSL100, and used it to generate queries. Generally, we observed a much longer time in query generation, and DSL100 frequently reached the attempt threshold before producing valid queries. This further resulted in lower quality of merged queries, and accordingly, worse detection results, as shown in Table 5. We also attempted to remove another 10% of random APIs from the DSL and used the remaining for query generation. It turned out that many previously successful queries could no longer be generated. This highlights the importance of performing a practical DSL subsetting to retain core features.
6.4.3. FP Elimination
During the iterative query generation, MoCQ treats those cases outside the generation dataset as false positives and uses them to refine the queries (§ 4.4.1). To understand the impact of such false positive signals, we experimented by disabling them as a comparison and showed the results as W/OFP in Table 5. Naturally, more false positives were introduced in the final reports when analyzing the projects in the testing dataset, even though the true positives remained the same. This in turn decreased the average F1 score to 0.3.
6.4.4. Generalization
Query generalization is important in the design of MoCQ. It enables MoCQ to capture diverse vulnerability patterns beyond the given input vulnerability example. Many new patterns could only be discovered with generalization. When we disabled it from MoCQ (W/OGen), the F1 score significantly dropped to 0.29. Our manual analysis of some queries revealed many new vulnerability patterns that the full-fledged MoCQ previously identified now disappeared. It is reasonable because by explicitly instrumenting the LLM to generalize the patterns, the LLM can better capture structural variations across vulnerabilities.
6.4.5. Models
MoCQ is not limited to only Claude 3.7 Sonnet. We further experimented with another model, GPT-4o (OpenAI, 2024). The results on Joern are shown in Table 5. In our experiments, ChatGPT-4o has an average F1 score of 0.4, while full-fledged MoCQ with Claude 3.7 Sonnet has the best performance of 0.51. Our experience suggests that the Claude 3.7 Sonnet model has a better capability in coding tasks than ChatGPT-4o. Though we did not integrate other models due to our cost constraints, we believe MoCQ has wide applicability.
6.4.6. Static Tools
Besides Joern, we further experimented MoCQ with CodeQL. CodeQL is capable of analyzing JavaScript but not PHP. We thus evaluated it using the same setup for the three types of JavaScript vulnerabilities in our dataset. From our experience, CodeQL queries are often considered harder due to their higher DSL complexity, so the results on two types of vulnerabilities were worse than MoCQ with Joern. It is interesting to observe that MoCQ with CodeQL achieved a better F1 score (0.72) on JavaScript prototype pollution.
Vul. Type MoCQ DSL100 W/OFP W/OGen GPT-4o CodeQL
PHP SQLi 0.65 0.38 0.54 0.37 0.44 -
PHP XSS 0.52 0.23 0.32 0.25 0.43 -
PHP Type. 0.25 0.15 0.19 0.23 0.31 -
PHP Deser. 0.59 0.41 0.34 0.28 0.50 -
JS Proto. 0.69 0.54 0.33 0.28 0.54 0.72
JS Cmd. 0.37 0.29 0.18 0.32 0.22 0.35
JS XSS 0.48 0.33 0.23 0.32 0.35 0.38
Average 0.51 0.33 0.30 0.29 0.40 0.48
6.5. DSL Subset
In this section, we evaluate and characterize the DSL subset extracted by MoCQ.
[…]
Number of Queries.Number of QueriesNumber of Queries. The query examples would have an impact on the DSL subset. We thus provided the Claude 3.7 Sonnet with progressively more query examples to understand how the amount of provided queries could impact the size core DSL subset. We conducted the experiments for Joern and CodeQL and presented the results in 6(a). Initially, more queries enriched the resulting core DSL subset, and the subset became gradually stable with 12 to 16 query examples. Ultimately, the subsetting technique could remove approximately half of the available APIs—48% for Joern and 55% for CodeQL—and significantly reduce the exploration space for query generation. Our earlier evaluation already demonstrated that the preserved subset was sufficiently expressive for vulnerability detection. While we do not claim this subset to be optimal or minimal, our manual analysis suggests that some preserved APIs could still be pruned. Nonetheless, the current design is effective and useful in practice.
Different choices of query examples might impact the resulting DSL subset. To measure this, we randomly selected 16 queries from a pool of 40 to construct the core DSL subset. We repeated this three times. The Venn diagram in 6(b) demonstrates that the three different query choices still resulted in a large overlap (93.8%). There could be some divergence across choices due to the variability in query examples. This suggests that the majority of the necessary features are consistent. We also measured the impact of using language-specific queries on the DSL subset in Appendix C.
[…]
Details of Removed Features.Details of Removed FeaturesDetails of Removed Features. We randomly sampled 50 removed APIs to explain why they were removed. We made a few key observations. First, the majority (45 out of 50) of removed DSL features in Joern could be equivalently expressed using the retained ones. For example, the whereNot() condition in Joern can be rewritten as a valid .where(not). This is particularly interesting, as DSL features are typically introduced to simplify usage and improve expressiveness (Hermans et al., 2009)—yet here we find that some offer redundant functionality that can be composed of other constructs. Two removed features are debugging interfaces and textual processing operations, such as printing. They are not necessary for production-environment queries. The other three are language-specific operations—those required only for specific programming languages. For example, address operation in Joern is applicable to binary code and not required for JavaScript or PHP. Similar findings were observed in CodeQL, though it has relatively more language-specific operations removed (7 out of 50).
(a) Query amount
(b) Query choices
Figure 6. Factors impacting the core DSL subset
@@ 7. Discussion
In this section, we discuss the current limitations and future directions.
[…]
Vulnerability Examples.Vulnerability ExamplesVulnerability Examples. At present, MoCQ requires a small set of true positive vulnerability samples, which helps the LLM understand the vulnerabilities and also serves as test cases to validate the queries. For our current evaluation, we randomly selected the vulnerability examples. A more careful selection of the vulnerability examples (e.g., cover diverse vulnerability perspectives) or providing more examples would advance the performance of MoCQ. Besides, we believe negative cases can be helpful to improve the precision of generated queries. We currently assume the reports outside the known true positives as the negative signals and wish the queries could exclude them. One can also additionally provide negative cases, including well-annotated true negatives and false negatives. For example, patched vulnerabilities can be a good source of true negatives.
[…]
Other Use Scenarios.Other Use ScenariosOther Use Scenarios. We present a generic pipeline for generating detection queries, which enables multiple practical use scenarios. For example, MoCQ can be used to enhance existing expert-crafted queries by automatically extending them with additional detection capabilities. We have showcased this for JavaScript prototype pollution. In addition, MoCQ can serve as a foundation for building a digital twin of an existing (black-box) static analysis tool. In this setting, feedback—such as true/false positives or partial results—from a closed-source tool can guide MoCQ in generating queries that approximate its behaviors. This is especially useful in industrial contexts where high-performing yet opaque tools lack extensibility or transparency. By mimicking such tools, MoCQ enables more flexible, explainable analysis for smooth integration into modern development workflows.
[…]
Model Fine-tuning.Model Fine-tuningModel Fine-tuning. We currently leverage in-context learning to generate detection queries in DSL. Our DSL specification mitigates the errors triggered during the generation process but MoCQ may still require many iterations to correct errors and refine the queries. Fine-tuning pretrained models could be an effective way to make MoCQ even more powerful. Our pipeline can then serve as a foundation to generate synthetic query data for resolving issues like lack of training data in fine-tuning. Our symbolic query validator can also be used as the reward function in reinforcement learning-based model training.
[…]
Ethical Considerations.Ethical ConsiderationsEthical Considerations. We strive to adhere to the best practices of research ethics throughout this project. We employed LLMs to generate vulnerability detection patterns in a controlled, local environment, ensuring that the generated outputs did not pose any risk to public systems or external stakeholders. Any newly discovered vulnerabilities were responsibly disclosed to the corresponding application developers. We are also actively engaging with the maintainers of expert-crafted vulnerability patterns to share our insights and improve existing queries.
@@ 8. Related Work
Learning-based Detection.Learning-based DetectionLearning-based Detection. Various transformer-based methods have been applied for vulnerability detection, including encoder-only  (Fen, 2020),I encoder-decoder (Wang et al., 2021, 2023a), and decoder-only (Touvron et al., 2023) architectures. Researchers have also adapted various fine-tuning strategies, such as domain-specific pretraining (Ding et al., 2023), instruct-tuning (Mao et al., 2024; Chang et al., 2024), supervised fine-tuning (Zhu et al., 2024; Wang et al., 2024b), parameter-efficient fine-tuning (Mao et al., 2024) to improve models’ performance. Complementary to fine-tuning approaches, prompt engineering using variations like few-shot, chain-of-thought, and progressive prompting strategies (Nong et al., 2025; Fu et al., 2023; Zhang et al., 2024; Liu et al., 2024b) have achieved promising results. However, multiple studies (Ullah et al., 2024) have identified limitations like response inconsistency and diminished efficacy when addressing under-represented vulnerabilities or large codebases. Compared with these works, MoCQ has the additional advantage of symbolic methods and is highly scalable.
[…]
Neuro-symbolic Program Analysis.Neuro-symbolic Program AnalysisNeuro-symbolic Program Analysis. Generally, neuro-symbolic program analysis combines a neural and a symbolic component. IRIS (Li et al., 2025) and Artemis (Ji et al., 2025) leverage LLMs to infer project-specific taint specifications for Java and PHP applications, respectively. They rely on signatures and comments to decide whether the functions are sources or sinks. IRIS further uses an LLM to eliminate false positives by providing the slices. There is also work that leverages neural methods to validate vulnerabilities, e.g., WAP (Medeiros et al., 2011). Furthermore, LLMSA (Wang et al., 2024a) uses an LLM to decompose the vulnerability analysis problem into several simple syntactic or semantic properties upon smaller code snippets, which are handled by the symbolic analysis rules. LLift (Li et al., 2024) uses an LLM to analyze the post constraints for use-before-initialization bugs within the Linux kernel. MoCQ distinguishes itself from these by targeting automated query generation from scratch, and we believe MoCQ can complement them to reduce engineering demands.
[…]
Cloned Vulnerability Detection.Cloned Vulnerability DetectionCloned Vulnerability Detection. Software developers might borrow code (thus vulnerabilities in it) from others. Multiple work uses the signature of a known vulnerability to detect such vulnerable code clones in other codebases. VUDDY (Kim et al., 2017) leverages a function-level granularity signature after code abstraction and normalization. It can only find extract clones without modifications (type-1) and renamed clones with identifier changes (type-2). TRACER (Kang et al., 2022b) represents the signature as an inter-procedural data-flow trace by collecting the operations in the trace; HiddenCPG (Wi et al., 2022) constructs a graph structure and RecurSan (Shi et al., 2024) calculates a symbolic expression for existing vulnerabilities. MoCQ is different by extracting and generalizing vulnerability patterns for a class of vulnerabilities instead of focusing on specific vulnerability instances.
[…]
General Static Analysis.General Static AnalysisGeneral Static Analysis. Lots of tools specifically target the analysis of PHP and JavaScript applications—the primary testing target of this paper—including TChecker (Luo et al., 2022), RIPS (Dahse and Holz, 2014), Pixy (Jovanovic et al., 2006), and ODGen (Li et al., 2022a). Unlike query-based systems, these tools typically rely on hard-coded analysis logic and do not support customizable or extensible analyses through user-defined queries. Realizing such customization on them requires additional implementation considerations and engineering. More generally, there is a line of efforts analyzing other programming languages such as C/C++ (Sui and Xue, 2016), Java (Vallée-Rai et al., 2010; Lam et al., 2011), Rust (Li et al., 2021b, 2022b), etc.
@@ 9. Conclusion
This work presented MoCQ, a novel neuro-symbolic approach that combines the complementary advantages of LLMs and static analysis to achieve fully automated vulnerability detection. With a DSL subsetting technique and trace-driven symbolic validation, MoCQ automatically extracts vulnerability patterns using an LLM. These patterns are then integrated with query-based static analysis to achieve high scalability in analyzing large codebases. Our evaluation demonstrates that MoCQ could achieve comparable detection performance as expert-crafted queries. It significantly reduces the required engineering efforts, lowers the barrier to building tailored vulnerability detectors, and makes automated security analysis more accessible and adaptable. We believe that our novel neuro-symbolic approach represents a promising new direction for general software security.
Appendix A Details of Preliminary Study
We conducted a preliminary study evaluating ChatGPT-4o’s capability to generate queries for five types of vulnerabilities among the OWASP Top 10 (OWASP Foundation, 2021). To invoke the LLM, we experimented with three different prompting strategies: (1) zero-shot, where the model is directly asked to generate a query for a vulnerability type without any query example; (2) chain-of-thought (CoT), which prompts the model to break down its reasoning step by step; and (3) few-shot, where the model is provided with two query examples. Each time, we also provided a true positive vulnerability example. The prompt templates are present in Table 9.
We evaluated the queries generated by ChatGPT-4o based on two key criteria: (1) execution success (Exe.), which checks whether a query runs without runtime errors, and (2) result correctness (Res.), which verifies if the query successfully identifies the intended vulnerability example. We did not assess whether the queries generalize to additional vulnerabilities beyond the input case. To reduce randomness, each experiment was repeated ten times.
Table 6 reports the success rates across ten trials for each vulnerability type. To our surprise, among all cases, only one query was successfully generated to retrieve a JavaScript command injection under the few-shot setting. Most of the queries raised errors or exceptions and could not be executed successfully. Compared to other settings, the few-shot approach offered limited guidance on DSL syntax and structure but was generally inadequate for producing valid or effective queries. In most cases, the generated queries failed to compile due to syntax or structural issues. Only a few compiled successfully, and among those, just one retrieved the intended vulnerability. This suggests that while few-shot examples provide partial insights, they are insufficient for generating functional queries in real-world scenarios. An iterative approach incorporating feedback or contextual information could further improve outcomes.
These issues are non-trivial to address, even for human experts. Two authors manually reviewed and attempted to fix the broken queries. Repairing syntax and structural errors alone often required several hours per query, underscoring both the complexity of the query languages and the steep learning curve involved.
Table 6. Evaluation of LLM-generated queries. Exe. and Res. denote execution success and result correctness, respectively.
Zero-shot CoT Few-shot
Vul. Type % Exe. % Res. % Exe. % Res. % Exe. % Res.
Joern PHP SQLi 0 0 0 0 30 0
PHP XSS 0 0 0 0 10 0
PHP Type. 0 0 0 0 0 0
JS Proto. 0 0 0 0 0 0
JS Cmd. 0 0 0 0 50 10
CodeQL JS Proto. 0 0 0 0 50 0
JS Cmd. 0 0 0 0 20 0
Appendix B Vulnerability Examples for High F1 Scores
In Table 7, we report the number of vulnerability examples required to reach such a stably high F1 score. We consider an F1 score high and stable when it no longer improves by more than 5% compared to the previous step in Figure 5. Note that each project may contribute a different number of examples. Despite this variation, the results indicate that fewer than 20 vulnerability examples are typically sufficient across different vulnerability types.
Table 7. Minimal number of examples for high F1 scores.
PHP SQLi PHP XSS PHP Type. PHP Deser. JS Proto. JS Cmd. JS XSS
15 18 5 13 14 14 6
Appendix C DSL Subset with Language-specific Queries
We assessed the impact of language-specific queries on the DSL subset. To explore this, we conducted another experiment on Joern comparing the core DSL subsets on PHP and JavaScript. We used two sets of 16 queries: one comprising general queries and PHP-specific queries, and the other comprising general queries and JavaScript-specific queries. Figure 7 shows that the two subsets overlapped by 95.11%. This demonstrates the target languages of the queries do not have a huge impact on obtaining the core subset. This also reflects that language-specific operations are not prevalent. We did not evaluate this in CodeQL because it does not support PHP.
Figure 7. Impact of the query languages on the DSL subset.
Appendix D Query Merging and Efficiency
The query merging stage in MoCQ could improve the analysis efficiency. We thus designed a baseline version by sequentially running each per-example query and accumulating the query execution time compared to the merged version. Table 8 shows the analysis time for each vulnerability type on the testing dataset. As we can see, the merging component (W/ M.) could finish the analysis several times faster than the unmerged one (W/O M.), demonstrating the necessity of query merging.
Table 8. Analysis time (in hours) with or without the merging component on the testing dataset.
PHP SQLi PHP XSS PHP Type. PHP Deser. JS Proto. JS Cmd. JS XSS
W/ M. 4.2h 4.9h 8.2h 1.5h 2.3h 0.4h 1.1h 0.7h
W/O M. 24+h 24+h 6.2h 5.4h 2.8h 3.4h 2.7h
Appendix E Prompt Design
We detail our prompt templates used in this project in Table 9.
Table 9. Prompt templates for various tasks in this paper.
Task Prompt
Preliminary Study (Appendix A) Zero-shot You are a security expert in analyzing PHP and JavaScript application vulnerabilities. You are proficient in tool-name. Given the following code snippet of a vulnerability-type, please generate a tool-name query that could be used to detect the given vulnerability-type vulnerability.
Chain-of-thought You are a security expert in analyzing PHP and JavaScript application vulnerabilities. You are proficient in tool-name. Given the following code snippet of a vulnerability-type, please generate a tool-name query that could be used to detect the given vulnerability-type vulnerability. Please provide step-by-step reasoning outlining how you arrive at the final query, including how you identify user-controlled data and its flow into dangerous methods.
Few-shot You are a security expert in analyzing PHP and JavaScript application vulnerabilities. You are proficient in tool-name. Given the following code snippet of a vulnerability-type, please generate a tool-name query that could be used to detect the given vulnerability-type vulnerability. We provide the following two query examples for your reference.
DSL Extraction (§ 4.1) Grammar You are an expert on the tool-name. You are given the documentation of tool-name. Your task is to analyze the documentation below to summarize the grammar of the domain-specific query language for the queries of tool-name. Please summarize the language’s grammar in the Backus-Naur Form.
Semantics You are an expert of the tool-name. You are given the source code implementations of tool-name. Your task is to analyze the implementations to summarize the functionalities or behaviors of the domain-specific APIs listed below.
Query Generation (§ 4.2) Generation You are an expert in tool-name. You will write a vulnerability detection query to detect vulnerability-type using the domain-specific language (DSL): {DSL syntax and semantics}. To do this, follow a step-by-step reasoning process to break down the detection task into smaller subtasks. For each step, generate a corresponding DSL query fragment. Finally, combine the fragments into a complete query. Here is a vulnerability example in vulnerability-type and a few query examples.
Refinement You are an expert on tool-name. Your previously generated query for vulnerability-type produced unexpected errors: syntax errors / execution exceptions / semantic errors. Please refer to the detailed error messages and our execution state to revise the query.
False-positives You are an expert on tool-name. Your previously generated query for vulnerability-type could successfully retrieve the input vulnerability example. However, it produced a few false positives: {false positive cases}, and the execution states concerning the false positives are: {execution states}. Please refer to that to revise the query.
Generalization You are an expert on tool-name. Your previously generated query for vulnerability-type could successfully retrieve the input vulnerability example. However, it is too specific to the input vulnerability example and is potentially overfitting for the pattern: {overfitting pattern}. Please refer to the program language specifications to generalize the query to also search other venerability variants beyond the input example.
LLM-based Detection (§ 6.3.2) Single-file analysis You are a security expert in analyzing PHP and JavaScript application vulnerabilities. Here are a few vulnerability examples in vulnerability-type. Please help me evaluate if the code snippet below contains any vulnerabilities in vulnerability-type. Here is the code to analyze: {code-in-a-file}.

### v3

@@ 5.5. RQ4: Ablation Study
We conduct a comprehensive ablation study to measure the individual contributions of the techniques on the testing dataset. We use the F1 score ($2\cdot\frac{Prec.\cdot Rec.}{Prec.+Rec.}$) as our evaluation metric, as it captures both precision and recall, and is well-suited for imbalanced data where non-vulnerable code is much more than vulnerable code. Due to space constraints, we summarize the main results and omit those for query merging and model temperature, though both of which were evaluated.
@@ 5.5.1. Vulnerability Examples
To understand how many vulnerability examples are needed for MoCQ to produce good detection results, starting from vulnerability examples in two projects, we gradually include more projects as inputs for MoCQ to generate queries. We then apply the queries to the testing dataset and compute the F1 scores. This incremental setup helps us understand the impact of additional projects on the detection effectiveness. We conduct our experiments at the project granularity rather than the vulnerability granularity due to similar data leakage concerns (discussed in § 5.1).
Our results show that F1 scores generally increase as the number of input projects grows, though the stabilization point varies by vulnerability type. For instance, JavaScript XSS queries reach stable high F1 scores with around four projects, while JavaScript prototype pollution requires closer to ten projects to stabilize. Overall, using about ten projects as input is sufficient for MoCQ to generate high-quality queries across most vulnerability types. This corresponds to roughly 5-18 vulnerability examples, depending on the type.
5.5.2. FP Elimination
During the iterative query generation, MoCQ treats those cases outside the generation dataset as false positives and uses them to refine the queries (§ 3.4.1). To understand the impact of such false positive signals, we experimented by disabling them as a comparison and showed the results as $W/O_{FP}$ in Table 4. Naturally, more false positives were introduced in the final reports when analyzing the projects in the testing dataset, even though the true positives remained the same. This in turn decreased the average F1 score to 0.29.
5.5.3. Generalization
Query generalization is important in the design of MoCQ. It enables MoCQ to capture diverse vulnerability patterns beyond the given input vulnerability example. Many new patterns could only be discovered with generalization. When we disabled it from MoCQ ($W/O_{Gen}$), the F1 score significantly dropped to 0.29. Our manual analysis of some queries revealed many new vulnerability patterns that the full-fledged MoCQ previously identified now disappeared. It is reasonable because by explicitly instrumenting the LLM to generalize the patterns, the LLM can better capture structural variations across vulnerabilities.
5.5.4. Models
MoCQ is not limited to only Claude 3.7 Sonnet. We further experimented with another model, GPT-4o (OpenAI, 2024). The results on Joern are shown in Table 4. In our experiments, GPT-4o has an average F1 score of 0.37, while full-fledged MoCQ with Claude 3.7 Sonnet has the best performance of 0.49. Our experience suggests that the Claude 3.7 Sonnet model has a better capability in coding tasks than GPT-4o. Though we did not integrate other models due to our cost constraints, we believe MoCQ has wide applicability.
5.5.5. Static Tools
Besides Joern, we experimented MoCQ with CodeQL. CodeQL is not capable of analyzing PHP. We thus evaluated it using the same setup for other vulnerabilities in our dataset. From our experience, CodeQL queries are often considered harder due to their higher DSL complexity, so the results on two types of vulnerabilities were worse than MoCQ with Joern. It is interesting to observe that MoCQ with CodeQL achieved a better F1 score (0.72) on JavaScript prototype pollution.
Vul. Type MoCQ W/OFP W/OGen GPT-4o CodeQL DSL100
PHP SQLi 0.65 0.54 0.37 0.44 - 0.38
PHP XSS 0.52 0.32 0.25 0.43 - 0.23
PHP type. 0.25 0.19 0.23 0.31 - 0.15
PHP deser. 0.59 0.34 0.28 0.50 - 0.41
JS proto. 0.69 0.33 0.28 0.54 0.72 0.54
JS Cmd. 0.37 0.18 0.32 0.22 0.35 0.29
JS XSS 0.48 0.23 0.32 0.35 0.38 0.33
C/C++ UAF 0.53 0.31 0.35 0.38 0.44 0.28
C/C++ heap ov. 0.46 0.27 0.22 0.30 0.39 0.23
Java dan. func 0.50 0.29 0.33 0.35 0.48 0.36
Java SQLi 0.55 0.26 0.29 0.32 0.40 0.30
Java API misuse 0.50 0.25 0.28 0.31 0.36 0.22
Average 0.51 0.29 0.29 0.37 0.44 0.31
@@ 5.5.6. DSL Subset
Joern exposes a total of 4,5914,5914 , 591 DSL APIs, while CodeQL exposes 12,93412,93412 , 934 DSL APIs. The subsetting process prunes 48% of Joern’s and 67% of CodeQL’s APIs, substantially reducing the search space for query generation while maintaining sufficient expressiveness for effective vulnerability detection.
The number of input queries influences the resulting DSL subset. We provided Claude 3.7 Sonnet with increasing numbers of query examples and observed that the subset size stabilized around 12 to 16 queries for both Joern and CodeQL. We omit the detailed diagram showing the stabilization curve due to space constraints. By sampling different sets but the same number of queries, we found a high overlap (93.8%) in the resulting DSL subsets, indicating robustness against variability in query examples.
We further manually analyzed 50 randomly sampled removed APIs; all of them were confirmed redundant, showing the reliability of LLMs in the DSL subsetting task. Most (45 out of 50) removed features had equivalent alternatives within the retained subset. For example, Joern’s whereNot() can be expressed as where(not). A few removed APIs were non-essential features such as debugging or printing operations, and some were language-specific APIs irrelevant to the target languages (e.g., address used only in binary analysis).
@@
To understand the effect of the subsetting technique, we developed a variant containing the full set of DSL, denoted as DSL100, and used it to generate queries. Generally, DSL100 resulted in much longer query generation times and frequently reached the attempt threshold before producing valid queries. This led to lower-quality merged queries and worse detection results, as shown in Table 4. We also experimented with removing an additional 10% of random APIs from the subset and found that many previously successful queries could no longer be generated. This highlights the importance of practical DSL subsetting to retain core, necessary features.
@@ 6. Discussion
Vulnerability Examples. We provided MoCQ with a small set of vulnerability examples to help it generate patterns for general vulnerability types. These examples serve as seed inputs from which MoCQ induces generalized patterns applicable to a broader class of similar vulnerabilities. They can be sourced from real-world data, such as CVE entries, security patches, or well-established vulnerability benchmarks. Importantly, the number of examples required is significantly smaller than those typically needed by deep learning-based techniques. Moreover, thanks to the code generation capabilities of modern LLMs, it is also feasible to generate high-quality synthetic vulnerability examples.
Extensibility. By design, MoCQ is language-agnostic and applicable to a wide range of languages and vulnerability types beyond those evaluated. It requires only a small number of vulnerability examples without requiring any changes to its design. Beyond Joern (Yamaguchi et al., 2014; Backes et al., 2017) and CodeQL (CodeQL, 2024a), MoCQ can be directly extended to frameworks such as Amazon CodeGuru (Services, 2025) and Datalog-based systems (Scholz et al., 2016). For other non-query-based static approaches, it may require specifying how the patterns are expressed and represented in these methods.
Other Use Scenarios. As a generic framework, MoCQ supports multiple practical use cases. For instance, it can enhance expert-developed queries by automatically extending them with additional detection logic, as demonstrated for JavaScript prototype pollution. MoCQ can also serve as a foundation for building a digital twin of closed-source static analyzers. In this setting, feedback such as true and false positives or partial matches can guide MoCQ to approximate the tool’s behavior. This is especially valuable in industrial contexts, where high-performing tools often lack transparency or extensibility. By mimicking such tools, MoCQ enables flexible, explainable analysis suitable for modern development workflows.
Ethical Considerations. We adhere to the best research ethics practices. We used LLMs in a local environment to generate vulnerability detection patterns. All newly discovered vulnerabilities and patterns were responsibly disclosed to the respective developers.
Threat to Validity. Our evaluation faces several threats to validity. First, our dataset is limited to 12 vulnerability types across four programming languages, and we did not experiment with others, though theoretically MoCQ is language-agnostic. Second, our comparison to expert-developed queries uses commit history as a proxy for development effort, which may not reflect actual time investment. Third, the manual collection of ground truth vulnerabilities may be incomplete, potentially affecting precision calculations.
@@ 7. Related Work
Learning-based Detection. Transformer-based models have been widely applied to vulnerability detection, including encoder-only (Fen, 2020), encoder-decoder (Wang et al., 2021, 2023a), and decoder-only (Touvron et al., 2023) architectures. To improve performance, researchers have explored various fine-tuning strategies, such as domain-specific pretraining (Ding et al., 2023), instruct-tuning (Mao et al., 2024; Chang et al., 2024), supervised fine-tuning (Zhu et al., 2024; Wang et al., 2024b), and parameter-efficient tuning (Mao et al., 2024). However, recent studies (Ullah et al., 2024) show that they require a large dataset for training and often present reduced effectiveness on under-represented vulnerabilities and large codebases. Compared to these approaches, MoCQ benefits from symbolic reasoning and achieves superior scalability, and requires only a small set of vulnerability examples.
Neuro-symbolic Program Analysis. Neuro-symbolic program analysis integrates neural and symbolic components. IRIS (Li et al., 2025) and Artemis (Ji et al., 2025) use LLMs to infer project-specific taint specifications for Java and PHP, respectively, based on signatures and comments. IRIS additionally applies LLMs to reduce false positives via slicing. Other works use neural components to assist symbolic validation, such as WAP (Medeiros et al., 2011), or decompose analysis into smaller property checks as in LLMSA (Wang et al., 2024a). LLift (Li et al., 2024) applies LLMs to reason about post-conditions in use-before-initialization bugs in the Linux kernel. MoCQ differs by aiming to automatically generate vulnerability queries from scratch, and can complement existing approaches by reducing manual engineering effort.
General Static Analysis. Static vulnerability detection includes a broad range of techniques beyond pattern-based approaches. Methods such as symbolic execution (Baldoni et al., 2018; Dudina and Belevantsev, 2017), model checking (Yang et al., 2006; Su et al., 2021), and formal verification (Hasan and Tahar, 2015) analyze program behavior against formal specifications or mathematical models, aiming to logically prove security properties or identify bugs. Other general techniques, such as dataflow analysis frameworks based on lattices (Cousot and Cousot, 1977), apply abstract interpretation to systematically reason about program behavior and detect potential vulnerabilities. Unlike pattern-based tools, these approaches do not depend on manually crafted vulnerability signatures. However, despite their precision and theoretical rigor, formal and dataflow-based methods often face significant scalability challenges when applied to large or complex codebases.
@@ 8. Conclusion
This work presents MoCQ, a novel neuro-symbolic approach that combines the complementary strengths of LLMs and static analysis to enable scalable, automated vulnerability detection. By leveraging a DSL subsetting technique and trace-driven symbolic validation, MoCQ automatically extracts vulnerability patterns using an LLM. MoCQ-generated queries achieved comparable detection performance to expert-developed queries with substantially reduced engineering effort. MoCQ lowers the barrier to developing custom vulnerability detectors and makes automated security analysis more accessible and adaptable. We believe this neuro-symbolic approach represents a promising direction for general software engineering.